# Nakai-Moishezon criteria and the toric Thomas-Yau conjecture

### v1

@@ Abstract
As an application, we prove that, on the mirror of a toric weak del Pezzo surface, if $\mathcal{L}$ defines a Bridgeland stable object in the Fukaya-Seidel category in a natural sense, then it is Hamiltonian isotopic to a special Lagrangian section in the class. The converse also holds on the mirror of $\operatorname{Bl}_{p}\mathbb{P}^{2}$. This is consistent with general conjectures due to Joyce.
@@ 1.1 Toric special Lagrangian sections and stability
In the present work we study a version of the Thomas-Yau conjecture in a toric case, namely, for a class of Lagrangian sections of certain Calabi-Yau Lagrangian fibrations, known as toric Landau-Ginzburg models of weak Fano type.
@@
Let $X$ be a toric weak Fano manifold, endowed with a fixed torus-invariant Kähler form $\omega\in[\omega_{0}]$. Let $L\to X$ be a holomorphic line bundle on $X$, endowed with a torus-invariant Hermitian metric $h$. By the classical work of Leung, Yau and Zaslow [28] (generalised to toric compact Kähler manifolds, as recalled e.g. in [5] and in [14], Section 9), one can construct a corresponding non-compact Lagrangian submanifold of a fixed bounded domain $\mathcal{U}_{q}$ in $(\mathbb{C}^{*})^{n}$,
@@
This is a section of a natural Lagrangian torus fibration $\mathcal{Y}_{q}\supset\mathcal{U}_{q}\to\Delta^{0}$ corresponding to $\omega$, where $\Delta^{0}$ denotes the interior of the momentum polytope. The Lagrangian section $\mathcal{L}(L,h)$ is known as the Strominger-Yau-Zaslow (SYZ) transform of $(L,h)$. Under this correspondence, changes in the metric $h$ preserve the Hamiltonian isotopy class of $\mathcal{L}$.
@@
Let $X$ be a toric weak Fano manifold with a fixed torus-invariant Kähler form $\omega_{0}$ and a holomorphic line bundle $L$. Suppose $L^{\vee}$ is ample. Fix $k>0$ sufficiently large. Let $\mathcal{L}=\mathcal{L}(L^{\otimes k},h)\subset\mathcal{U}_{q_{k}}$ be a SYZ Lagrangian section in the mirror to $(X,k\omega_{0})$. Then, as $V$ ranges through toric subvarieties of $X$, there exist objects
@@ 1.2 Relation to Bridgeland stability on surfaces
Let $X$ be a toric weak del Pezzo surface. Suppose the fixed line bundle $L^{\vee}$ is semiample. Fix $k>0$ sufficiently large. Consider a SYZ Lagrangian section $\mathcal{L}=\mathcal{L}(L^{\otimes k},h)$. We say that the Lagrangian
@@
Definition 1.5 is consistent. Taking the shift ~$\tilde{\mathcal{L}}=\mathcal{L}[1]$ in the category $D\!\operatorname{FS}((\mathcal{Y}_{q_{k}},W(k\omega_{0}))$ is represented by the same Lagrangian section $\mathcal{L}$, endowed with the new grading obtained by shifting the lifted Lagrangian phase of $\mathcal{L}$ by $\pi$, see [29], footnote 1 and Remark 2.6. On the other hand, by the cohomological characterisation of the heart $\operatorname{Coh}^{\sharp}(X)$ recalled in Section 6, the shifted line bundle $L^{\otimes k}[1]$ defines an object of $\operatorname{Coh}^{\sharp}(X)$.
@@
Proposition 1.8.
In the situation of Theorem 1.2 and Definition 1.5, when $X$ is a toric weak del Pezzo surface, if $\tilde{\mathcal{L}}$ is Bridgeland stable, then it is Hamiltionian isotopic to a (shifted) special SYZ Lagrangian section. In the special case when $X=\operatorname{Bl}_{p}\mathbb{P}^{2}$ the converse also holds.
The proof of Proposition is given in Section 6.1.
@@
The proof of Proposition 1.8 is based on showing that, when $X$ is a toric surface (not necessarily weak del Pezzo), Bridgeland stability of $L^{\otimes k}[1]$ with respect to $\sigma_{k[\omega_{0}]}$ implies the existence of a dHYM connection on $L^{\otimes k}$, see Proposition 6.1. That the converse holds on $X=\operatorname{Bl}_{p}\mathbb{P}^{2}$ is a result of Collins and Shi [11]. We note that this is compatible with the wall-and-chamber decompositions for dHYM connections studied by Khalid and Sjöström Dyrefelt [26].
@@
Collins and Shi [11] show an example of a Kähler class $[\omega_{0}]$ and a line bundle $L$ on $X=\operatorname{Bl}_{p}\mathbb{P}^{2}$ which is stable with respect to stability condition $(\operatorname{Coh}^{\sharp}(X),Z_{[\omega_{0}]})$, but for which the dHYM equation is not solvable, implying that $\tilde{\mathcal{L}}$ is not Hamiltonian isotopic to a (shifted) special SYZ Lagrangian section. This $L^{\vee}$ is seminegative, and so does not satisfy the assumptions of Definition 1.5 and Proposition 1.8, although probably the example can be modified.
@@
Acknowledgements. I am very grateful to Tristan Collins, Carlo Scarpa, Nicolò Sibilla and Richard Thomas for important comments and suggestions.
2 Background and main results
@@
We will always assume that $X$ is toric. In this case, it is enough to test (2.3) on toric submanifolds $V\subset X$. Furthermore, we will always work with torus-invariant representatives of $[\omega_{0}]$. Then a solution $\alpha\in[\alpha_{0}]$ is necessarily torus-invariant, by uniqueness.
@@ 2.2 Toric weak Fano manifolds
Suppose $X$ is a toric weak Fano manifold. Then, for all sufficiently large, fixed $k>0$, there exist classes
@@
where $V$ ranges over all toric submanifolds, such that the dHYM positivity condition (2.3) for $(X,[\omega_{0}],c_{1}(L^{\vee}))$ is equivalent to the phase inequalitites for periods
@@ 2. $(ii)$
    see Remark 5.5. We will apply this in Section 6 in order to describe a relation with Bridgeland stability conditions when $X$ is a toric weak del Pezzo surface, and in particular to prove Proposition 1.8.
@@ 3. $(iii)$
    The proof of Theorem 2.5 also shows that, in principle, the scale factor $k>0$ could be controlled quantitatively in terms of Givental’s $I$-function $I(q,-z)$ for $X$ and so in terms of genus 00 Gromov-Witten invariants (we recall the definition of $I(q,-z)$ in Section 5.1).
@@ 2.4 Higher rank
for fixed line bundles $L_{i}\to X$. Let $E:=E_{1}$. If there exists $h\in\mathcal{H}^{+}(E)$ solving the higher rank dHYM equation with $B$-field on $E$, i.e. (2.5), then, for all sufficiently large $k>0$, we have
$\arg\int_{\Gamma}e^{-W(k(\omega_{0}-\operatorname{i}\beta_{0}))}\Omega_{0}-\pi<\arg\left(\int_{\Gamma_{V}}e^{-W(k(\omega_{0}-\operatorname{i}\beta_{0}))}\Omega_{0}\right)-\pi<\arg\int_{\Gamma}e^{-W(k(\omega_{0}-\operatorname{i}\beta_{0}))}\Omega_{0},$
@@
$\arg\int_{[\mathcal{L}]}e^{-W(k(\omega_{0}-\operatorname{i}\beta_{0}))}\Omega_{0}<\arg\left(\int_{[\mathcal{L}_{V}]}e^{-W(k(\omega_{0}-\operatorname{i}\beta_{0}))}\Omega_{0}\right)-\pi<\arg\int_{[\mathcal{L}]}e^{-W(k(\omega_{0}-\operatorname{i}\beta_{0}))}\Omega_{0},$
@@
The definition of $\sigma_{k[\omega_{0}],k[\beta_{0}]}$ is recalled in Section 8.3. We will see that Definition 2.18 is consistent (Section 8.3) and has the correct central charge (see (8.3)). We also have an analogue of Proposition 1.8.
@@
In the situation of Theorem 2.16, fix $k>0$ sufficiently large, and suppose that the $k[\beta_{0}]$-twisted slopes of $L^{\otimes k}_{i}$ are nonpositive, $\mu_{k[\omega_{0}],k[\beta_{0}]}(L^{\otimes k}_{i})\leq 0$, $i=1,2$. If $\mathcal{L}_{V}$ or $\mathcal{L}_{1}$ violate the phase inequalities for periods appearing in Corollary 2.17, then $\tilde{\mathcal{L}}$ is Bridgeland unstable, destabilised by $\mathcal{L}_{V}$ or $\mathcal{L}_{1}$.
@@ 2.5 General toric manifolds
Let $X$ be a projective toric manifold with a fixed Kähler class $[\omega_{0}]$ and a line bundle $L\to X$. For all sufficiently large $k>0$, there exist complex cycles $\Gamma_{L}$, $\Gamma_{L,V}$, as $V$ ranges through toric submanifolds of $X$, representing classes
@@ 5.2.1 Divisors
Thus, by (5.4) and (5.5), the phase inequality (3.2) (which is equivalent to dHYM positivity (2.3)), namely
$\arg\left((-1)^{\operatorname{codim}V}\int_{V}e^{-\operatorname{i}\omega_{0}}\operatorname{ch}(L)\right)<\arg\left(-\int_{X}e^{-\operatorname{i}\omega_{0}}\operatorname{ch}(L)\right),$ (5.8)
can be expressed as
@@
$\int_{V}e^{-\operatorname{i}\omega_{0}}\operatorname{ch}(L)=k^{-n+1}k^{-1}_{V}\int_{X}e^{-\operatorname{i}k\omega_{0}}\operatorname{ch}(\mathcal{S}(kL,k_{V}V))+O(k^{-1}_{V}).$

### v2

@@ Abstract
As an application, we prove that, on the mirror of a toric weak del Pezzo surface, if $\mathcal{L}$ defines a Bridgeland stable object in the Fukaya-Seidel category in a natural sense, then it is Hamiltonian isotopic to a special Lagrangian section in the class. The converse also holds for the mirror of $\operatorname{Bl}_{p}\mathbb{P}^{2}$ and $\operatorname{Bl}_{p,q}\mathbb{P}^{2}$, and always holds assuming a conjecture of Arcara and Miles. When $\mathcal{L}$ is Bridgeland unstable, we obtain a morphism from $\mathcal{L}$ to a weak solution of the special Lagrangian equation with phase angle satisfying a minimality condition. These results are consistent with general conjectures due to Joyce.
@@ 1.1 Toric special Lagrangian sections and stability
Here we study a version of the Thomas-Yau conjecture in a toric case, namely, for a class of Lagrangian sections of certain Calabi-Yau Lagrangian fibrations, known as toric Landau-Ginzburg models of weak Fano type.
@@
Let $X$ be a toric weak Fano manifold, endowed with a fixed torus-invariant Kähler form $\omega_{0}$. Let $L\to X$ be a holomorphic line bundle on $X$, endowed with a torus-invariant Hermitian metric $h$. By the classical work of Leung, Yau and Zaslow [29] (generalised to toric compact Kähler manifolds, as recalled e.g. in [5] and in [14], Section 9), one can construct a corresponding non-compact Lagrangian submanifold of a fixed bounded domain $\mathcal{U}_{q}$ in $(\mathbb{C}^{*})^{n}$,
@@
This is a section of a natural Lagrangian torus fibration $\mathcal{Y}_{q}\supset\mathcal{U}_{q}\to\Delta^{0}$ corresponding to $\omega_{0}$, where $\Delta^{0}$ denotes the interior of the momentum polytope. The Lagrangian section $\mathcal{L}(L,h)$ is known as the Strominger-Yau-Zaslow (SYZ) transform of $(L,h)$. Under this correspondence, changes in the metric $h$ preserve the Hamiltonian isotopy class of $\mathcal{L}$.
@@
Let $X$ be a toric weak Fano manifold with a fixed torus-invariant Kähler form $\omega_{0}$ and a holomorphic line bundle $L$, with ample inverse $L^{\vee}$. Fix $k>0$ sufficiently large. Let $\mathcal{L}=\mathcal{L}(L^{\otimes k},h)\subset\mathcal{U}_{q_{k}}$ be a SYZ Lagrangian section in the mirror to $(X,k\omega_{0})$. Suppose $[\omega_{0}]$ is generic in the sense that it does not lie in the union of finitely many proper analytic subvarieties of $H^{1,1}(X,\mathbb{R})$ determined by $c_{1}(L)$.
Then, as $V$ ranges through toric subvarieties of $X$, there exist objects
@@
Remark 1.4.
The genericity condition on $[\omega_{0}]$ is explicit, see Remark 2.6 $(v)$.
@@ 1.2 Relation to Bridgeland stability on surfaces
Let $X$ be a toric weak del Pezzo surface. Suppose the fixed line bundle $L^{\vee}$ is positive. Fix $k>0$ sufficiently large. Let $\mathcal{L}=\mathcal{L}(L^{\otimes k},h)\subset\mathcal{U}_{q_{k}}$ be a SYZ Lagrangian section in the mirror to $(X,k\omega_{0})$. We say that the Lagrangian
@@
Definition 1.6 is consistent. Taking the shift ~$\tilde{\mathcal{L}}=\mathcal{L}[1]$ in the category $D\!\operatorname{FS}(\mathcal{Y}_{q_{k}},W(k\omega_{0}))$ is represented by the same Lagrangian section $\mathcal{L}$, endowed with the new grading obtained by shifting the lifted Lagrangian phase of $\mathcal{L}$ by $\pi$, see [30], footnote 1 and Remark 2.6. On the other hand, by the cohomological characterisation of the heart $\operatorname{Coh}^{\sharp}(X)$ recalled in Section 6, the shifted line bundle $L^{\otimes k}[1]$ defines an object of $\operatorname{Coh}^{\sharp}(X)$.
@@
Theorem 1.9.
Suppose we are in the situation of Theorem 1.2 and Definition 1.6, with $X$ is a toric weak del Pezzo surface. Then, if $\tilde{\mathcal{L}}$ is Bridgeland stable, it is Hamiltionian isotopic to a (shifted) special SYZ Lagrangian section. The converse also holds, conditionally on a conjecture of Arcara-Miles (see Conjecture 6.1), which is known when $X=\operatorname{Bl}_{p}\mathbb{P}^{2}$ or $X=\operatorname{Bl}_{p,q}\mathbb{P}^{2}$.
The proof is given in Section 6.0.1.
@@
The proof of Theorem 1.9 is based on showing that, under the assumptions of Theorem 1.2, when $X$ is a toric surface (not necessarily weak del Pezzo), Bridgeland stability of $L^{\otimes k}[1]$ with respect to $\sigma_{k[\omega_{0}]}$ for $k\gg 1$ implies the existence of a dHYM connection on $L^{\otimes k}$, i.e. a solution of (2.1), see Proposition 6.2. Conversely, using an argument of Collins and Shi [11], we will see that the dHYM positivity of $L^{\otimes k}$ with respect to $k[\omega_{0}]$ implies the Bridgeland stability of $L^{\otimes k}[1]$ with respect to $\sigma_{k[\omega_{0}]}$ for all $k\geq 1$, if the conjecture of Arcara-Miles holds. The conjecture was proved for $X=\operatorname{Bl}_{p}\mathbb{P}^{2}$ in [3] and recently for $X=\operatorname{Bl}_{p,q}\mathbb{P}^{2}$ in [31].
We note that Theorem 1.9 is compatible with the wall-and-chamber decompositions for dHYM connections studied by Khalid and Sjöström Dyrefelt [27].
@@
Collins and Shi [11] show an example of a Kähler class $[\omega_{0}]$ and a line bundle $L$ on $X=\operatorname{Bl}_{p}\mathbb{P}^{2}$ which is stable with respect to stability condition $(\operatorname{Coh}^{\sharp}(X),Z_{[\omega_{0}]})$, but for which the dHYM equation (2.1) is not solvable, implying that $\tilde{\mathcal{L}}$ is not Hamiltonian isotopic to a (shifted) special SYZ Lagrangian section. Their $L^{\vee}$ is negative, and so does not satisfy the assumptions of Definition 1.6 and Theorem 1.9, although probably the example can be modified.
@@
1.2.1 The unstable case
The works of Joyce [25] and Li [30] also contain precise expectations concerning the case of objects $\mathcal{L}$ in the Fukaya category which are unstable (i.e. not semistable) with respect to the (conjectural) Bridgeland stability condition relevant for the existence of special Lagrangians. These relate to versions of the Harder-Narashiman filtration for $\mathcal{L}$.
In our very special situation, when $X$ is a toric weak del Pezzo surface, we can make some progress on the unstable case by relying on the work of Datar-Mete-Song [15] concerning weak solutions of the dHYM equation of surfaces. These are briefly recalled in Section 6.1.
Let $X$ be a toric weak del Pezzo surface. Suppose the fixed ample line bundle $L^{\vee}$ on $X$ does not support a smooth solution of the dHYM equation (2.1), so that $\mathcal{L}=\mathcal{L}(L,h)\subset\mathcal{U}_{q}$ cannot be Hamiltinian isotopic to a smooth special SYZ Lagrangian section. In this case, following the results of [15], there exists a minimal slope $\mathbb{R}$-divisor $D$ in $X$, which attains the minimal angle $\theta_{min}\in(0,\pi)$ determined by
$\cot\theta_{min}:=\sup_{D}\left\{\frac{(c_{1}(L^{\vee})-D)^{2}-[\omega_{0}]^{2}}{2(c_{1}(L^{\vee})-D)\cdot[\omega_{0}]},\,(c_{1}(L^{\vee})-D)\cdot[\omega_{0}]>0\right\},$
where the supremum is taken over effective $\mathbb{R}$-divisors.
Note that, if a line bundle $\hat{L}$ satisfies $c_{1}(\hat{L}^{\vee})\cdot[\omega_{0}]>0$ and supports a smooth solution of the dHYM equation (2.1), then the phase angle $\varphi(X,\hat{L},[\omega_{0}])\in(0,\pi)$ of the mirror special SYZ Lagrangian section is given by
$\cot\varphi(X,\hat{L},[\omega_{0}])=\frac{(c_{1}(\hat{L}^{\vee}))^{2}-[\omega_{0}]^{2}}{2c_{1}(\hat{L}^{\vee})\cdot[\omega_{0}]}.$
Theorem 1.12.
Suppose we are in the situation of Theorem 1.2 and Definition 1.6, with $X$ is a toric weak del Pezzo surface. Let the Lagrangian $\tilde{\mathcal{L}}$ be Bridgeland unstable (i.e. not semistable) with respect to $\sigma_{k[\omega_{0}]}$. Suppose that the corresponding minimal slope divisor $D$ is in fact a $\mathbb{Q}$-divisor. Then, assuming the Arcara-Miles conjecture 6.1 (known for $X=\operatorname{Bl}_{p}\mathbb{P}^{2}$ or $X=\operatorname{Bl}_{p,q}\mathbb{P}^{2}$), possibly after increasing $k>0$, there is a (shifted) SYZ Lagrangian section ~$\tilde{\mathcal{L}}_{D}$, with a morphism
~~$\tilde{\mathcal{L}}\to\tilde{\mathcal{L}}_{D}$
in $D\!\operatorname{FS}(\mathcal{Y}_{q_{k}},W(k\omega_{0}))$, such that ~$\tilde{\mathcal{L}}_{D}$ is Hamiltonian isotopic to a (shifted) SYZ Lagrangian section in the sense of currents, which is a weak solution of the special Lagrangian equation with phase angle $\theta_{min}$.
The technical details of the statement and its proof are given in Section 6.1.1 (see also Example 6.6, following [15], Theorem 1.13).
@@ 1.2.1 The unstable case
Acknowledgements. I am grateful to Vamsi Pingali, Carlo Scarpa, Nicolò Sibilla, Peng Zhou and especially to Tristan Collins and Richard Thomas for important comments and suggestions.
2 Background and main result
@@
We will always assume that $X$ is toric. In this case, it is enough to test (2.3) on toric submanifolds $V\subset X$. Furthermore, we will always work with a torus-invariant representative $\omega_{0}$. Then a solution $\alpha\in[\alpha_{0}]$ is necessarily torus-invariant, by uniqueness.
@@ 2.2 Toric weak Fano manifolds
Let $X$ be a toric weak Fano manifold with a fixed Kähler class $\omega_{0}$ and a holomorphic line bundle $L$. For all sufficiently large, fixed $k>0$, there exist classes
@@
where $V$ ranges over all toric submanifolds of $X$, such that, if $[\omega_{0}]$ is generic in the sense that it does not lie in the union of finitely many proper analytic subvarieties of $H^{1,1}(X,\mathbb{R})$ determined by $c_{1}(L)$, then the dHYM positivity condition (2.3) for $(X,[\omega_{0}],c_{1}(L^{\vee}))$ is equivalent to the phase inequalitites for periods
@@ 2. $(ii)$
    see Remark 5.5. We will apply this in Section 6 in order to describe a relation with Bridgeland stability conditions when $X$ is a toric weak del Pezzo surface, and in particular to prove Theorem 1.9.
@@ 3. $(iii)$
    The proof of Theorem 2.5 shows that, in principle, the scale factor $k>0$ could be controlled quantitatively in terms of Givental’s $I$-function $I(q,-z)$ for $X$ and so in terms of genus 00 Gromov-Witten invariants, virtually enumerating rational curves contained in $X$ (we recall the definition of $I(q,-z)$ in Section 5.1). The role of holomorphic curves in the Thomas-Yau conjecture has been emphasised by Li [30].
@@ 4. $(iv)$
  5. $(v)$
    The proof of Theorem 2.5 shows that the genericity condition required on $[\omega_{0}]$ is precisely the absence of strict semi-stabilisers $V$ for the Nakai-Moishezon criterion (2.3), i.e. $V\subset X$ for which the left hand side of (2.3) vanishes.
@@ 2.4 Higher rank
for fixed line bundles $L_{i}\to X$. Let $E:=E_{1}$.
If there exists $h\in\mathcal{H}^{+}(E)$ solving the higher rank dHYM equation with $B$-field (2.5) on $E$, then, for all sufficiently large $k>0$, if $[\omega_{0}]$ is generic in the sense that it does not lie in the union of finitely many proper analytic subvarieties of $H^{1,1}(X,\mathbb{R})$ determined by $c_{1}(L_{i})$, $i=1,2$, we have
$\arg\int_{\Gamma}e^{-W(k(\omega_{0}-\operatorname{i}\beta_{0}))}\Omega_{0}<\arg\left(\int_{\Gamma_{V}}e^{-W(k(\omega_{0}-\operatorname{i}\beta_{0}))}\Omega_{0}\right)<\arg\int_{\Gamma}e^{-W(k(\omega_{0}-\operatorname{i}\beta_{0}))}\Omega_{0}+\pi,$
@@
$\arg\int_{[\mathcal{L}]}e^{-W(k(\omega_{0}-\operatorname{i}\beta_{0}))}\Omega_{0}<\arg\left(\int_{[\mathcal{L}_{V}]}e^{-W(k(\omega_{0}-\operatorname{i}\beta_{0}))}\Omega_{0}\right)<\arg\int_{[\mathcal{L}]}e^{-W(k(\omega_{0}-\operatorname{i}\beta_{0}))}\Omega_{0}+\pi,$
@@
The definition of $\sigma_{k[\omega_{0}],k[\beta_{0}]}$ is recalled in Section 8.3. We will see that Definition 2.18 is consistent (Section 8.3) and has the correct central charge (see (8.3)). We also have an analogue of Theorem 1.9.
@@
In the situation of Theorem 2.16, fix $k>0$ sufficiently large, and suppose that the $k[\beta_{0}]$-twisted slopes of $L^{\otimes k}_{i}$ are nonpositive, $\mu_{k[\omega_{0}],k[\beta_{0}]}(L^{\otimes k}_{i})\leq 0$, $i=1,2$.
If $\mathcal{L}_{V}$ or $\mathcal{L}_{1}$ violate the phase inequalities for periods appearing in Corollary 2.17, then $\tilde{\mathcal{L}}$ is Bridgeland unstable, destabilised by $\mathcal{L}_{V}$ or $\mathcal{L}_{1}$.
@@ 2.5 General toric manifolds
Let $X$ be a projective toric manifold with a fixed Kähler class $[\omega_{0}]$ and a line bundle $L\to X$.
For all sufficiently large $k>0$, if $[\omega_{0}]$ is generic in the sense that it does not lie in the union of finitely many proper analytic subvarieties of $H^{1,1}(X,\mathbb{R})$ determined by $c_{1}(L)$, then there exist complex cycles $\Gamma_{L}$, $\Gamma_{L,V}$, as $V$ ranges through toric submanifolds of $X$, representing classes
@@ 5.2.1 Divisors
Then, the above argument, in particular (5.4) and (5.2.1), show that the phase inequality (3.2) (which is equivalent to dHYM positivity (2.3)), namely
$\arg\left((-1)^{\operatorname{codim}V}\int_{V}e^{-\operatorname{i}\omega_{0}}\operatorname{ch}(L)\right)<\arg\left(\int_{X}e^{-\operatorname{i}\omega_{0}}\operatorname{ch}(L)\right),$ (5.8)
implies
@@
Conversely, if we assume the inequality (5.9), then we get the inequality of leading order terms in the expansion for $k\gg k_{V}\gg 1$,
$\arg\left((-1)^{\operatorname{codim}V}\int_{V}e^{-\operatorname{i}\omega_{0}}\operatorname{ch}(L)\right)\leq\arg\left(\int_{X}e^{-\operatorname{i}\omega_{0}}\operatorname{ch}(L)\right).$
In general, the inequality may not be strict. However, this is ruled out if we suppose that $[\omega_{0}]$ is generic with respect to $V$ in the sense that it does not lie in the proper analytic subvariety of $H^{1,1}(X,\mathbb{R})$ determined by $c_{1}(L)$ and $V$ through the condition
$\arg\left((-1)^{\operatorname{codim}V}\int_{V}e^{-\operatorname{i}\omega_{0}}\operatorname{ch}(L)\right)=\arg\left(\int_{X}e^{-\operatorname{i}\omega_{0}}\operatorname{ch}(L)\right),$
which, under our assumptions, is precisely the condition that $V$ is strictly semi-stabilising in the Nakai-Moishezon criterion,
$\int_{V}\operatorname{Re}(\operatorname{i}\omega_{0}+c_{1}(L^{\vee}))^{\dim V}-\cot(\varphi)\operatorname{Im}(\operatorname{i}\omega_{0}+c_{1}(L^{\vee}))^{\dim V}=0.$ (5.10)
@@
Then the equivalence between the conditions (5.8) and (5.9) holds as long as $[\omega_{0}]$ is generic with respect to all divisors in the sense that it does not lie in the union of finitely many proper analytic subvarieties of $H^{1,1}(X,\mathbb{R})$ determined by $c_{1}(L)$ and the classes of all toric divisors.
@@
$\int_{V}e^{-\operatorname{i}\omega_{0}}\operatorname{ch}(L)=k^{-n+1}k^{-1}_{V}\int_{X}e^{-\operatorname{i}k\omega_{0}}\operatorname{ch}(\mathcal{S}(kL,k_{V}V))+O(k^{-1}_{V}),$
which follows from our computations above. Thus, the phase inequality (3.2) (which is equivalent to dHYM positivity (2.3)), namely
$\arg\left((-1)^{\operatorname{codim}V}\int_{V}e^{-\operatorname{i}\omega_{0}}\operatorname{ch}(L)\right)<\arg\left(\int_{X}e^{-\operatorname{i}\omega_{0}}\operatorname{ch}(L)\right),$
can be expressed as
$\arg\left((-1)^{\operatorname{codim}V}\int_{X}e^{-\operatorname{i}k\omega_{0}}\operatorname{ch}(\mathcal{S}(kL,k_{V}V))\right)<\arg\int_{X}e^{-\operatorname{i}k\omega_{0}}\operatorname{ch}(L^{\otimes k}),$
where we have a morphism $\mathcal{S}(kL,k_{V}V)[-\operatorname{codim}V]\to L^{\otimes k}$ in $D^{b}(X)$, provided that $[\omega_{0}]$ is generic with respect to all divisors in the sense that it does not lie in the union of finitely many proper analytic subvarieties of $H^{1,1}(X,\mathbb{R})$ determined by $c_{1}(L)$ and the classes of all toric divisors through the conditions (5.10) (as $V$ varies).
We will use this fact in the proof of Theorem 1.9, concerning Bridgeland stability conditions for Lagrangian sections, when $X$ is a toric weak del Pezzo surface.